# refrences.bib

### v1

@@ 6 Numerical example and application
Table 6: Updated Table 5 after applying Corollary 12.
$\{0.4\}$$\{0.4\}$varnothing $\varnothing$ $\{0.6\}$
$\varnothing$ $\{0.64\}$$\{0.6\}$$\{1\}$varnothing
$\{0.4\}$varnothing $\{0.6\}$$\{1\}$$\{0.3\}$
The situation of Table 6 and the reduced matrices $A^{+}$ and $A^{-}$are the same as having three rows (rows 1,4 and 5 in the main problem) and five columns 1-5. As it is shown in Table 6, $\mathscr{I}_{1}=\mathscr{I}_{5}=\{1,3\}$ and $\bigcap_{i\in\mathscr{I}_{1}}S_{i1}^{\prime}=S_{11}^{\prime}\cap S_{31}^{\prime}=\{0.4\}$. On the other hand, $L_{1}=0.4$ (see Table 3), and therefore Corollary 13 (Part (a)) implies that $x_{5}^{*}=L_{5}=0.3$. Also, column 5 can be removed from the problem. Similarly, since $\mathscr{I}_{3}=\mathscr{I}_{4}=\{2,3\}$, $\bigcap_{i\in\mathscr{I}_{3}}S_{i3}^{\prime}=\{0.6\}=\left\{U_{3}\right\}$, $\bigcap_{i\in\mathscr{I}_{4}}S_{i4}^{\prime}=\{1\}=\left\{U_{4}\right\}$ and $c_{3}\left(U_{3}-L_{3}\right)=0.93(0.6-0)=0.558<3.28=3.28(1-0)=c_{4}\left(U_{4}-L_{4}\right)$, from Part (b) of Corollary 13 , we can set $x_{4}^{*}=L_{4}=0$ and remove column 4 . By applying this simplification technique, for the remaining rows (rows 1, 4 and 5 in the main problem), sets $\mathscr{J}_{1}=\{1,2,5\}$, $\mathscr{J}_{4}=\{2,3,4\}$ and $\mathscr{J}_{5}=\{1,3,4,5\}$ are updated as $\mathscr{J}_{1}=\{1,2\}$, $\mathscr{J}_{4}=\{2,3\}$ and $\mathscr{J}_{5}=\{1,3\}$, respectively. So, we have $\prod_{i\in\mathscr{I}-\{2,3,6,7,8,9,10\}}\left|\mathscr{J}_{i}\right|=2\times 2\times 2=8$ and reduced matrices $A^{+}$ and $A^{-}$, vector $b$ and Table 6 are updated as follows:
@@
Table 7: Updated Table 6 after applying Corollary 13.
$\{0.4\}$$\{0.4\}$varnothing
@@
Figure 1 contains the solution tree generated by the branch and bound approach, which is based on the concept of modified functions. This method begins by choosing $e(1)$ from $\bar{\mathscr{J}}_{1}(e)=\mathscr{J}_{1}=\{1,2\}$. Hence, each $e\in\bar{E}$ must satisfy either $e(1)=1$ or $e(1)=2$. This yields two branches denoted by nodes 1 and 2 in Figure 1 . For node 1, by considering $e_{1}\in\bar{E}$ and $e_{1}(1)=1$, we have $e_{1}=[1]$, $\mathscr{I}_{1}\left(e_{1}\right)=\{1\}$ and $\mathscr{I}_{2}\left(e_{1}\right)=\mathscr{I}_{3}\left(e_{1}\right)=\varnothing$. So, from (15) we obtain $x\left(e_{1}\right)_{1}=\min\left\{\bigcap_{i\in\mathscr{I}_{1}\left(e_{1}\right)}S_{i1}^{\prime}\right\}=\min S_{11}^{\prime}=0.4$, $x\left(e_{1}\right)_{2}=L_{2}=0.4$ and $x\left(e_{1}\right)_{3}=L_{3}=0$ with $z_{1}=c_{1}x\left(e_{1}\right)_{1}+c_{2}x\left(e_{1}\right)_{2}+c_{3}x\left(e_{1}\right)_{3}=1(0.4)+0.35(0.4)+0.93(0)=0.54$. Similarly, by selecting $e(1)=2$ and considering another admissible function $e_{2}\in\bar{E}$ such that $e_{2}(1)=2$, it follows that $e_{2}=[2]$, $x\left(e_{2}\right)=[0.4,0.4,0]$ and $z_{2}=0.54$. At this time, we have no reason to exclude nodes 1 and 2 from consideration; therefore, both nodes should be further investigated. By selecting node 1 (where $e_{1}(1)=1$ ), for the second row (row 4 in the main problem), we have $\mathscr{J}_{4}=\{2,3\}$. Since $\mathscr{I}_{2}\left(e_{1},2\right)=\mathscr{I}_{3}\left(e_{1},2\right)=\varnothing$, then $\mathscr{J}_{2}\left(e_{1}\right)=\{2,3\}$ (Definition 6). Also, since $\mathscr{J}_{2}\left(e_{1}\right)\cap\left\{e_{1}(1)\right\}=\varnothing$, we have $\bar{\mathscr{J}}_{2}\left(e_{1}\right)=\mathscr{J}_{2}\left(e_{1}\right)=\{2,3\}$ (Definition 9). This implies that for each $e\in\bar{E}$ such that $e(1)=1$ , we must have either $e(2)=2$ or $e(2)=3$. So, branching on node 1 yields nodes 3 and 4 in Figure 1, which lead us to two admissible functions $e_{1}$ and $e_{3}$, respectively. At node 3, $e_{1}(2)=2$ and therefore we obtain $e_{1}=[1,2]$, $x\left(e_{1}\right)=[0.4,0.64,0]$ (resulted by (15)) and $z_{1}=c_{1}x\left(e_{1}\right)_{1}+c_{2}x\left(e_{1}\right)_{2}+c_{3}x\left(e_{1}\right)_{3}=1(0.4)+0.35(0.64)+0.93(0)=0.624$. On the other hand, if we move along $e(2)=3$ to node 4 , similar reasoning shows that $e_{3}=[1,3]$, $x\left(e_{3}\right)=[0.4,0.4,0.6]$ and $z_{3}=1.098$ for node 4. Again, there is no reason to exclude nodes 3 and 4 from consideration, so we need to branch on one node. By using the jump-tracking technique, a branch is made at a node that has a lower bound on $z$. Here, node 3 has been chosen.
@@
At node 3, $e_{1}(1)=1$ and $e_{1}(2)=2$. For the third row (row 5 in the main problem), we have $\mathscr{J}_{5}=\{1,3\}$. Since $\mathscr{I}_{1}\left(e_{1},3\right)=\{1\}$, $S_{31}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{1}\left(e_{1},3\right)}S_{k1}^{\prime}\right)=S_{31}^{\prime}\cap S_{11}^{\prime}=\{0.4\}$ and $\mathscr{I}_{3}\left(e_{1},3\right)=\varnothing$, then $\mathscr{J}_{3}\left(e_{1}\right)=\{1,3\}\quad$ (Definition 6). However, $\mathscr{J}_{3}\left(e_{1}\right)\cap\left\{e_{1}(1),e_{1}(2)\right\}=\mathscr{J}_{3}\left(e_{1}\right)\cap\{1,2\}=\{1\}$, which means $\bar{\mathscr{J}}_{3}\left(e_{1}\right)=\{1\}$ (Definition 9). This implies that $e_{1}(3)=1$ is the only selection for constructing a modified function. So, any choice associated with node 3 must satisfy $e_{1}(3)=1$, which yields one branch from node 3, namely, node 5 in Figure 1. By the same reasoning as before, for this node, we have $e_{1}=[1,2,1]$, $x\left(e_{1}\right)=[0.4,0.64,0]$ and $z_{1}=0.624$. Node 5 corresponds to the complete branch $e_{1}=[1,2,1]$ and $x\left(e_{1}\right)=[0.4,0.64,0]$ may be viewed as a candidate solution with $z_{1}=0.624$.
@@
Because the objective value for node 4 ($z=1.098$) cannot be lower than 0.624, this node can be eliminated from further consideration. Hence, we now branch on node 2. At node 2, $e_{2}(1)=2$. So, at the second row (row 4 in the main problem), we have $\mathscr{J}_{4}=\{2,3\}$. Since $\mathscr{I}_{2}\left(e_{2},2\right)=\{1\}$, $S_{22}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{2}\left(e_{2},2\right)}S_{k2}^{\prime}\right)=S_{22}^{\prime}\cap S_{12}^{\prime}=\{0.6\}\cap\{0.4\}=\varnothing$ and $\mathscr{I}_{3}\left(e_{2},2\right)=\varnothing$, then $\mathscr{J}_{2}\left(e_{2}\right)=\{3\}$ (Definition 6). Also, $\mathscr{J}_{2}\left(e_{2}\right)\cap\left\{e_{2}(1)\right\}=\mathscr{J}_{2}\left(e_{2}\right)\cap\{2\}=\varnothing$, which means $\bar{\mathscr{J}}_{2}\left(e_{2}\right)=\mathscr{J}_{2}\left(e_{2}\right)=\{3\}$ (Definition 9). This implies that $e_{2}(2)=3$ is the only selection for constructing a modified function. So, any modified function associated with node 2 must satisfy $e_{2}(2)=3$, that yields one branch from node 2 . Correspondingly, we create node 6 in Figure 1, where $e_{2}=[2,3]$, $x\left(e_{2}\right)=[0.4,0.4,0.6]$ and $z_{2}=1.098$. Thus, since $z_{2}>z_{1}$, this node is eliminated from consideration. As a result, the branch and bound method finds the optimal variables $x_{1}^{*}=0.4$, $x_{2}^{*}=0.64$ and $x_{3}^{*}=0$ (associated with the solution $x\left(e_{1}\right)=[0.4,0.64,0]$ ) by checking only one complete modified function $e_{1}=[1,2,1]$.
@@
Step 6.
According to Steps 4 and 5, an optimal solution of the problem is obtained as
@@ Conclusion
In this paper, an algorithm was proposed to find a global optimal solution to linear problems that involve a wide class of bipolar fuzzy relation equations defined with continuous Archimedean t-norms (including both continuous strict and nilpotent t-norms). Several basic properties of continuous Archimedean bipolar equations were explored, and the feasible solution set of the problem was completely characterized by a finite number of compact sets. In addition to determining the feasibility of the problem, two necessary feasibility conditions were also presented. In addition, seven simplification techniques were proposed to facilitate the process of solution finding; four techniques for deleting redundant constraints and/or eliminating redundant columns by assigning a fixed value to their corresponding variables, and three additional techniques for reducing the domain search to a subset of the feasible region, which contains the optimal solutions. Furthermore, a branch-and-bound-based algorithm was presented that takes advantage of the concept of modified functions to accelerate the solution process by neglecting some candidate solutions that are not optimal and removing some branches that do not lead to an optimal solution.
@@
{tblr}
c—c—c t-norm & Function Parameter
Product $T_{P}(x,y)=xy$
Einstein Product $T_{EP}(x,y)=\frac{xy}{2-(x+y-xy)}$
Lukasiewicz $T_{L}(x,y)=max\{0,x+y-1\}$
@@
Table A2: Additive generator $f_{\varphi}$ and pseudoinverse $f_{\varphi}^{(-1)}$ discussed in Lemma 2 for the t-norms stated in Table A1.
@@
Table A3: Value $u$ discussed in Theorems 1 and 2 for the t-norms stated in Table A1.
@@
Product $b/a$ Lukasiewicz $1+b-a$
Einstein Product $((2-a)b)/(a+b-ab)$ Frank $\log_{s}\left(1+\left[\left(s^{b}-1\right)(s-1)/\left(s^{a}-1\right)\right]\right)$
@@
    Suppose that $e\in E$ and $\mathscr{I}_{j_{0}}(e)\neq\varnothing$ for some $j_{0}\in\mathscr{J}$. Also, without loss of generality, let $i_{0}=\max\mathscr{I}_{j_{0}}(e)$. So, according to Definition 6 , we have $\mathscr{I}_{j_{0}}(e)=\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\cup\left\{i_{0}\right\}$ and $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{J_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. Thus, $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. To prove the converse statement, since $e(i)\in\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), it is sufficient to show that if $e\left(i_{0}\right)=j_{0}$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$, then $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. However, since $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, the statement $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$ implies $\mathscr{I}_{j_{0}}(e)\neq\varnothing$. Hence, from the assumption, we have $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. Therefore, since $i_{0}\in\mathscr{I}_{j_{0}}(e)$ (because $\left.e\left(i_{0}\right)=j_{0}\right)$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, it is concluded that $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. ∎
  • The proof of Theorem 3.
@@
    Let $x\in\bigcup_{e\in E}S(e)$. So, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. Hence, according to (11), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}\left(e_{0}\right)\neq\varnothing$ ) or $x_{j}\in I_{j}$ (if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$ ). But, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 4), from $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ it is obtained again $x_{j}\in I_{j}$. Consequently, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$. On the other hand, from Definition 6 we have $e_{0}(i)=j_{i}\in\mathscr{J}_{i}\left(e_{0}\right)\subseteq\mathscr{J}_{i}$ $(\forall i\in\mathscr{I})$, which implies $\mathscr{I}_{j_{i}}\left(e_{0}\right)\neq\varnothing$, and therefore from (11) we have $x_{j_{i}}\in\bigcap_{k\in\mathscr{I}_{j_{i}}\left(e_{0}\right)}S_{kj_{i}}^{\prime}\subseteq S_{ij_{i}}^{\prime}$. Hence, $x_{j_{i}}\in S_{ij_{i}}^{\prime}$, $\forall i\in\mathscr{I}$. Now, Lemma 5 requires that $x\in S\left(A^{+},A^{-},b\right)$. Conversely, let $x\in S\left(A^{+},A^{-},b\right),~{}\mathscr{I}_{j}(x)=\left\{i\in\mathscr{I}:x_{j}\in S_{ij}^{\prime}\right\}$ and $\mathscr{J}_{i}(x)=\left\{j\in\mathscr{J}:x_{j}\in S_{ij}^{\prime}\right\}$. So, for each $i\in\mathscr{I}_{j}(x)$ and each $j\in\mathscr{J}_{i}(x)$, we have $S_{ij}^{\prime}\neq\varnothing$ that means $j\in\mathscr{J}_{i}$ (Definition 5). Also, Lemma 5 implies that $\mathscr{J}_{i}(x)\neq\varnothing$, $\forall i\in\mathscr{I}$. Without loss of generality, let $j_{i}=\min\mathscr{J}_{i}(x)$ and $e_{0}(i)=j_{i}$, $\forall i\in\mathscr{I}$. Therefore, $e_{0}$ is a function on $\mathscr{I}$ such that
@@
    Consequently, by Corollary 6 and (12) - (14) we have $e_{0}\in E$. Also, since $x\in S\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$ (Lemma 5). Particularly, $x_{j}\in I_{j}$, if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$. This fact together with (11), (13) and (14) imply $x\in S\left(e_{0}\right)$. ∎
  • The proof of Lemma 6.

### v2

@@ 6 Numerical example and application
Table 6: Updated Table 5 after applying Corollary 5.14.
$\{0.4\}$ $\{0.4\}$ $\varnothing$ $\varnothing$ $\{0.6\}$
$\varnothing$ $\{0.64\}$ $\{0.6\}$ $\{1\}$ $\varnothing$
$\{0.4\}$ $\varnothing$ $\{0.6\}$ $\{1\}$ $\{0.3\}$
The situation of Table 6 and the reduced matrices $A^{+}$ and $A^{-}$are the same as having three rows (rows 1,4 and 5 in the main problem) and five columns 1-5. As it is shown in Table 6, $\mathscr{I}_{1}=\mathscr{I}_{5}=\{1,3\}$ and $\bigcap_{i\in\mathscr{I}_{1}}S_{i1}^{\prime}=S_{11}^{\prime}\cap S_{31}^{\prime}=\{0.4\}$. On the other hand, $L_{1}=0.4$ (see Table 3), and therefore Corollary 5.16 (Part (a)) implies that $x_{5}^{*}=L_{5}=0.3$. Also, column 5 can be removed from the problem. Similarly, since $\mathscr{I}_{3}=\mathscr{I}_{4}=\{2,3\}$, $\bigcap_{i\in\mathscr{I}_{3}}S_{i3}^{\prime}=\{0.6\}=\left\{U_{3}\right\}$, $\bigcap_{i\in\mathscr{I}_{4}}S_{i4}^{\prime}=\{1\}=\left\{U_{4}\right\}$ and $c_{3}\left(U_{3}-L_{3}\right)=0.93(0.6-0)=0.558<3.28=3.28(1-0)=c_{4}\left(U_{4}-L_{4}\right)$, from Part (b) of Corollary 5.16 , we can set $x_{4}^{*}=L_{4}=0$ and remove column 4 . By applying this simplification technique, for the remaining rows (rows 1, 4 and 5 in the main problem), sets $\mathscr{J}_{1}=\{1,2,5\}$, $\mathscr{J}_{4}=\{2,3,4\}$ and $\mathscr{J}_{5}=\{1,3,4,5\}$ are updated as $\mathscr{J}_{1}=\{1,2\}$, $\mathscr{J}_{4}=\{2,3\}$ and $\mathscr{J}_{5}=\{1,3\}$, respectively. So, we have $\prod_{i\in\mathscr{I}-\{2,3,6,7,8,9,10\}}\left|\mathscr{J}_{i}\right|=2\times 2\times 2=8$ and reduced matrices $A^{+}$ and $A^{-}$, vector $b$ and Table 6 are updated as follows:
@@
Table 7: Updated Table 6 after applying Corollary 5.16.
$\{0.4\}$ $\{0.4\}$ $\varnothing$
@@
Figure 2 contains the solution tree generated by the branch and bound approach, which is based on the concept of modified functions. This method begins by choosing $e(1)$ from $\bar{\mathscr{J}}_{1}(e)=\mathscr{J}_{1}=\{1,2\}$. Hence, each $e\in\bar{E}$ must satisfy either $e(1)=1$ or $e(1)=2$. This yields two branches denoted by nodes 1 and 2 in Figure 2 . For node 1, by considering $e_{1}\in\bar{E}$ and $e_{1}(1)=1$, we have $e_{1}=[1]$, $\mathscr{I}_{1}\left(e_{1}\right)=\{1\}$ and $\mathscr{I}_{2}\left(e_{1}\right)=\mathscr{I}_{3}\left(e_{1}\right)=\varnothing$. So, from (12) we obtain $x\left(e_{1}\right)_{1}=\min\left\{\bigcap_{i\in\mathscr{I}_{1}\left(e_{1}\right)}S_{i1}^{\prime}\right\}=\min S_{11}^{\prime}=0.4$, $x\left(e_{1}\right)_{2}=L_{2}=0.4$ and $x\left(e_{1}\right)_{3}=L_{3}=0$ with $z_{1}=c_{1}x\left(e_{1}\right)_{1}+c_{2}x\left(e_{1}\right)_{2}+c_{3}x\left(e_{1}\right)_{3}=1(0.4)+0.35(0.4)+0.93(0)=0.54$. Similarly, by selecting $e(1)=2$ and considering another admissible function $e_{2}\in\bar{E}$ such that $e_{2}(1)=2$, it follows that $e_{2}=[2]$, $x\left(e_{2}\right)=[0.4,0.4,0]$ and $z_{2}=0.54$. At this time, we have no reason to exclude nodes 1 and 2 from consideration; therefore, both nodes should be further investigated. By selecting node 1 (where $e_{1}(1)=1$ ), for the second row (row 4 in the main problem), we have $\mathscr{J}_{4}=\{2,3\}$. Since $\mathscr{I}_{2}\left(e_{1},2\right)=\mathscr{I}_{3}\left(e_{1},2\right)=\varnothing$, then $\mathscr{J}_{2}\left(e_{1}\right)=\{2,3\}$ (Definition 3.6). Also, since $\mathscr{J}_{2}\left(e_{1}\right)\cap\left\{e_{1}(1)\right\}=\varnothing$, we have $\bar{\mathscr{J}}_{2}\left(e_{1}\right)=\mathscr{J}_{2}\left(e_{1}\right)=\{2,3\}$ (Definition 5.17). This implies that for each $e\in\bar{E}$ such that $e(1)=1$ , we must have either $e(2)=2$ or $e(2)=3$. So, branching on node 1 yields nodes 3 and 4 in Figure 2, which lead us to two admissible functions $e_{1}$ and $e_{3}$, respectively. At node 3, $e_{1}(2)=2$ and therefore we obtain $e_{1}=[1,2]$, $x\left(e_{1}\right)=[0.4,0.64,0]$ (resulted by (12)) and $z_{1}=c_{1}x\left(e_{1}\right)_{1}+c_{2}x\left(e_{1}\right)_{2}+c_{3}x\left(e_{1}\right)_{3}=1(0.4)+0.35(0.64)+0.93(0)=0.624$. On the other hand, if we move along $e(2)=3$ to node 4 , similar reasoning shows that $e_{3}=[1,3]$, $x\left(e_{3}\right)=[0.4,0.4,0.6]$ and $z_{3}=1.098$ for node 4. Again, there is no reason to exclude nodes 3 and 4 from consideration, so we need to branch on one node. By using the jump-tracking technique, a branch is made at a node that has a lower bound on $z$. Here, node 3 has been chosen.
@@
At node 3, $e_{1}(1)=1$ and $e_{1}(2)=2$. For the third row (row 5 in the main problem), we have $\mathscr{J}_{5}=\{1,3\}$. Since $\mathscr{I}_{1}\left(e_{1},3\right)=\{1\}$, $S_{31}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{1}\left(e_{1},3\right)}S_{k1}^{\prime}\right)=S_{31}^{\prime}\cap S_{11}^{\prime}=\{0.4\}$ and $\mathscr{I}_{3}\left(e_{1},3\right)=\varnothing$, then $\mathscr{J}_{3}\left(e_{1}\right)=\{1,3\}\quad$ (Definition 3.6). However, $\mathscr{J}_{3}\left(e_{1}\right)\cap\left\{e_{1}(1),e_{1}(2)\right\}=\mathscr{J}_{3}\left(e_{1}\right)\cap\{1,2\}=\{1\}$, which means $\bar{\mathscr{J}}_{3}\left(e_{1}\right)=\{1\}$ (Definition 5.17). This implies that $e_{1}(3)=1$ is the only selection for constructing a modified function. So, any choice associated with node 3 must satisfy $e_{1}(3)=1$, which yields one branch from node 3, namely, node 5 in Figure 2. By the same reasoning as before, for this node, we have $e_{1}=[1,2,1]$, $x\left(e_{1}\right)=[0.4,0.64,0]$ and $z_{1}=0.624$. Node 5 corresponds to the complete branch $e_{1}=[1,2,1]$ and $x\left(e_{1}\right)=[0.4,0.64,0]$ may be viewed as a candidate solution with $z_{1}=0.624$.
@@
Because the objective value for node 4 ($z=1.098$) cannot be lower than 0.624, this node can be eliminated from further consideration. Hence, we now branch on node 2. At node 2, $e_{2}(1)=2$. So, at the second row (row 4 in the main problem), we have $\mathscr{J}_{4}=\{2,3\}$. Since $\mathscr{I}_{2}\left(e_{2},2\right)=\{1\}$, $S_{22}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{2}\left(e_{2},2\right)}S_{k2}^{\prime}\right)=S_{22}^{\prime}\cap S_{12}^{\prime}=\{0.6\}\cap\{0.4\}=\varnothing$ and $\mathscr{I}_{3}\left(e_{2},2\right)=\varnothing$, then $\mathscr{J}_{2}\left(e_{2}\right)=\{3\}$ (Definition 3.6). Also, $\mathscr{J}_{2}\left(e_{2}\right)\cap\left\{e_{2}(1)\right\}=\mathscr{J}_{2}\left(e_{2}\right)\cap\{2\}=\varnothing$, which means $\bar{\mathscr{J}}_{2}\left(e_{2}\right)=\mathscr{J}_{2}\left(e_{2}\right)=\{3\}$ (Definition 5.17). This implies that $e_{2}(2)=3$ is the only selection for constructing a modified function. So, any modified function associated with node 2 must satisfy $e_{2}(2)=3$, that yields one branch from node 2 . Correspondingly, we create node 6 in Figure 2, where $e_{2}=[2,3]$, $x\left(e_{2}\right)=[0.4,0.4,0.6]$ and $z_{2}=1.098$. Thus, since $z_{2}>z_{1}$, this node is eliminated from consideration. As a result, the branch and bound method finds the optimal variables $x_{1}^{*}=0.4$, $x_{2}^{*}=0.64$ and $x_{3}^{*}=0$ (associated with the solution $x\left(e_{1}\right)=[0.4,0.64,0]$ ) by checking only one complete modified function $e_{1}=[1,2,1]$.
@@
Step 6.7.
According to Steps 6.5 and 6.6, an optimal solution of the problem is obtained as
@@ Conclusion
This paper proposes an algorithm to find a global optimal solution to linear problems involving a wide class of bipolar fuzzy relation equations defined with continuous Archimedean t-norms (including both continuous strict and nilpotent t-norms). There were several basic properties of continuous Archimedean bipolar equations explored in this study, and a finite number of compact sets characterized the feasible solution set. Additionally, two necessary feasibility conditions were presented as part of the feasibility determination. As well, seven simplification techniques were introduced to facilitate the process of solution finding; four techniques for deleting redundant constraints and/or eliminating redundant columns by assigning a fixed value to their corresponding variables, and three additional techniques for reducing the domain search to a subset of the feasible region, which contains the optimal solutions. Then, a branch-and-bound-based algorithm was presented that takes advantage of the concept of modified functions to accelerate the solution process by neglecting some candidate solutions that are not optimal and removing some branches that do not lead to an optimal solution. It is worth pointing out that it is a future study to consider the assumption of left-hand continuity throughout the paper, and the comparison result is well worth assessing. Despite this, since the complexity of solving the problem increases exponentially with the problem size, it is always a challenge to find other simplification rules or methods that can reach the optimal solution.
@@
t-norm Function Parameter
Product $T_{P}(x,y)=xy$ -
Einstein Product $T_{EP}(x,y)=\frac{xy}{2-(x+y-xy)}$ -
Lukasiewicz $T_{L}(x,y)=max\{0,x+y-1\}$ -
@@
Table A2: Additive generator $f_{\varphi}$ and pseudoinverse $f_{\varphi}^{(-1)}$ discussed in Lemma 2.3 for the t-norms stated in Table A1.
@@
Table A3: Value $u$ discussed in Theorems 2.5 and 2.6 for the t-norms stated in Table A1.
@@
Product $b/a$ Lukasiewicz 1+ba1+b-a
Einstein Product ((2a)b)/(a+bab)((2-a)b)/(a+b-ab) Frank $\log_{s}\left(1+\left[\left(s^{b}-1\right)(s-1)/\left(s^{a}-1\right)\right]\right)$
@@
Suppose that $e\in E$ and $\mathscr{I}_{j_{0}}(e)\neq\varnothing$ for some $j_{0}\in\mathscr{J}$. Also, without loss of generality, let $i_{0}=\max\mathscr{I}_{j_{0}}(e)$. So, according to Definition 3.6 , we have $\mathscr{I}_{j_{0}}(e)=\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\cup\left\{i_{0}\right\}$ and $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{J_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. Thus, $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. To prove the converse statement, since $e(i)\in\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), it is sufficient to show that if $e\left(i_{0}\right)=j_{0}$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$, then $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. However, since $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, the statement $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$ implies $\mathscr{I}_{j_{0}}(e)\neq\varnothing$. Hence, from the assumption, we have $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. Therefore, since $i_{0}\in\mathscr{I}_{j_{0}}(e)$ (because $\left.e\left(i_{0}\right)=j_{0}\right)$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, it is concluded that $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. ∎
The proof of Theorem 3.10.
@@
Let $x\in\bigcup_{e\in E}S(e)$. So, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. Hence, according to (11), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}\left(e_{0}\right)\neq\varnothing$ ) or $x_{j}\in I_{j}$ (if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$ ). But, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 2.15), from $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ it is obtained again $x_{j}\in I_{j}$. Consequently, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$. On the other hand, from Definition 3.6 we have $e_{0}(i)=j_{i}\in\mathscr{J}_{i}\left(e_{0}\right)\subseteq\mathscr{J}_{i}$ $(\forall i\in\mathscr{I})$, which implies $\mathscr{I}_{j_{i}}\left(e_{0}\right)\neq\varnothing$, and therefore from (11) we have $x_{j_{i}}\in\bigcap_{k\in\mathscr{I}_{j_{i}}\left(e_{0}\right)}S_{kj_{i}}^{\prime}\subseteq S_{ij_{i}}^{\prime}$. Hence, $x_{j_{i}}\in S_{ij_{i}}^{\prime}$, $\forall i\in\mathscr{I}$. Now, Lemma 3.2 requires that $x\in S\left(A^{+},A^{-},b\right)$. Conversely, let $x\in S\left(A^{+},A^{-},b\right),~\mathscr{I}_{j}(x)=\left\{i\in\mathscr{I}:x_{j}\in S_{ij}^{\prime}\right\}$ and $\mathscr{J}_{i}(x)=\left\{j\in\mathscr{J}:x_{j}\in S_{ij}^{\prime}\right\}$. So, for each $i\in\mathscr{I}_{j}(x)$ and each $j\in\mathscr{J}_{i}(x)$, we have $S_{ij}^{\prime}\neq\varnothing$ that means $j\in\mathscr{J}_{i}$ (Definition 3.5). Also, Lemma 3.2 implies that $\mathscr{J}_{i}(x)\neq\varnothing$, $\forall i\in\mathscr{I}$. Without loss of generality, let $j_{i}=\min\mathscr{J}_{i}(x)$ and $e_{0}(i)=j_{i}$, $\forall i\in\mathscr{I}$. Therefore, $e_{0}$ is a function on $\mathscr{I}$ such that
@@
Consequently, by Corollary 3.8 and (12) - (14) we have $e_{0}\in E$. Also, since $x\in S\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$ (Lemma 3.2). Particularly, $x_{j}\in I_{j}$, if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$. This fact together with (11), (13) and (14) imply $x\in S\left(e_{0}\right)$. ∎
The proof of Lemma 5.1.